# SimGen: Simulator-conditioned Driving Scene Generation

### v1

@@ Abstract
Controllable synthetic data generation can substantially lower the annotation cost of training data in autonomous driving research and development. Prior works use diffusion models to generate driving images conditioned on the 3D object layout. However, those models are trained on small-scale datasets like nuScenes, which lack appearance and layout diversity. Moreover, the trained models can only generate images based on the real-world layout data from the validation set of the same dataset, where overfitting happens. In this work, we introduce a simulator-conditioned scene generation framework called SimGen that can learn to generate diverse driving scenes by mixing data from the simulator and the real world. It uses a novel cascade diffusion pipeline to address challenging sim-to-real gaps and multi-condition conflicts. A driving video dataset DIVA is collected to enhance the generative diversity of SimGen, which contains over 147.5 hours of real-world driving videos from 73 locations worldwide and simulated driving data from the MetaDrive simulator. SimGen achieves superior generation quality and diversity while preserving controllability based on the text prompt and the layout pulled from a simulator. We further demonstrate the improvements brought by SimGen for synthetic data augmentation on the BEV detection and segmentation task and showcase its capability in safety-critical data generation. Code, data, and models will be made available at https://metadriverse.github.io/simgen/.
Figure 1: We propose SimGen, a controllable scene generation paradigm through the simulator-conditioned generative model. It learns from real-world and simulated data and then can generate diverse driving scenes based on the simulator’s control conditions and text prompt.
This work is done during YZ’s visit to the University of California, Los Angeles.Corresponding author.
@@ 1 Introduction
A high-quality and diverse training data corpus is crucial for autonomous driving research and development. However, it is costly and laborious to annotate the data. Synthetic data generation is a promising alternative to harvest annotated training data, which brings realistic images and notable performance improvements across tasks like object detection [9] and semantic segmentation [73]. Besides the realism of the generated images, there are two necessary conditions for a practical synthetic data generator for autonomous driving: 1) Appearance diversity, which ensures the synthetic data can cover a spectrum of weather, environmental, and geographical conditions. 2) Layout diversity, namely the distribution of objects should cover different traffic scenarios, including safety-critical situations that are rare to collect in the real world.
@@
We construct a new dataset called DIVA to increase the appearance and layout diversity of the training data. DIVA comprises two parts: the web data and the synthesized data from the simulator. On the one hand, web data covers a worldwide range of geography, weather, scenes, and traffic elements, preserving the appearance diversity of a wide range of traffic participants. We design a data curation pipeline to collect and label YouTube driving videos. On the other hand, virtual driving videos with the traffic flow replayed from trajectory datasets or generated by a safety-critical scenario generator [78] are collected from a driving simulator [34]. In short, DIVA dataset blends real-world appearances and virtual layouts, consisting of 147.5 hours of Diverse In-the-wild and Virtual driving datA.
@@ 2.2 DIVA-Sim: Layout Diversity from the Simulator
Figure 2: Construction of DIVA dataset. DIVA-Real (left) comprises driving videos collected from YouTube. We apply a Vision Language Model to filter out noisy images via a checklist and utilize off-the-shelf models to annotate text, depth, and semantic labels. Meanwhile, DIVA-Sim (right) employs scene records and control policies in a simulator to create map elements and objects. It can generate digital twins of real-world data and safety-critical scenes. Then various kinds of sensors placed in the simulation produce multimodal images. Ren.:rendered; T.D.: top-down view. Numbers and letters indicate the sequence of processes.
@@ 3 SimGen Framework
Figure 3: Illustration of SimGen. SimGen processes text and scene record as inputs. The text is feature-encoded and utilized in the subsequent modules, whereas the scene record undergoes a simulator rendering into simulated depth and segmentation (SimCond) and extra conditions (ExtraCond). SimCond, coupled with the text features, is fed into the CondDiff module that converts SimCond into RealCond, representing real depth and segmentation. Eventually, the text features, SimCond, and ExtraCond are inputted into the ImgDiff module, where an Adapter merges multi-source conditions into a unified control condition and generates driving scene images.
@@ Appendix A Discussions
SimGen project page provide links to the YouTube videos (DIVA-Real_Video_Links) used in DIVA-Real, as well as digital twins of nuScenes dataset (DIVA-Sim_nuSc_Digital_Twins) and safety-critical video clips (DIVA-Sim_Safety-critical_Demo_Videos) included in DIVA-Sim.
@@
Q3. Broader impact. What are potential applications and future directions with the provided DIVA data and the SimGen model, for both academia and industry?
@@
Q4. Limitations. What are the issues with current designs, and corresponding preliminary solutions?
@@
However, as a pioneering work exploring how to introduce simulators into generative models for diverse driving scene generation, the primary focus of this work is the simulator-conditioned generalization ability across unseen driving scenarios rather than multi-view designs and computational overhead. Future work might include trying cross-frame attentions [18, 36, 70], faster sampling methods [44, 61, 82], and transferring our general method to more efficient diffusion models.
@@ B.1.2 Multimodal Annotation
Table 7: Location distribution of DIVA-Real.
Continent North America South America Europe Asian Africa
Proportion 56.9% 8.5% 16.9% 14.6% 3.1%
Table 8: Time period distribution of DIVA-Real.
Period Daytime Dawn Dusk Nighttime
Proportion 55.8% 16.3% 10.1% 17.8%
Table 9: Weather distribution of DIVA-Real.
Weather Normal Rainy Cloudy Foggy Snowy
Proportion 58.2% 1.0% 28.6% 2.1% 10.2%
@@ B.2.2 Layout Diversity Highlights
Table 10: Layout distribution of DIVA-Sim.
Behavior Forward Left Right Left Lane Right Lane Intersection U-Turn Stop
@@ B.4 License and Privacy Considerations
All the data is under the CC BY-NC-SA 4.0 license [^1] https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en. Other datasets (including nuScenes [51], Waymo Open [63], Metadrive [34]) inherit their own distribution licenses.
@@ C.2 Model Design
Realism-controllability trade-off. Apart from the discretization steps of the CondDiff solver, the critical hyperparameter for sim-to-real transformation is $t_{s}$, the starting time of the image synthesis process in the reverse SDE. We notice that with a fixed CondDiff model, there’s a trade-off between Realism and Controllability when choosing different $t_{s}$ values. Smaller $t_{s}$ values lead to fewer denoising steps, giving SimCond more control over image generation but potentially compromising realism. Generally, we find $t_{s}\in[0.4,0.65]$ to work well for the foreground, and we ultimately select $t_{s}$ as 0.5 for foregrounds. Then, we mask the foreground part, and there will be an additional $t_{s}=0.5$ for background generation.

### v2

@@ Abstract
Controllable synthetic data generation can substantially lower the annotation cost of training data. Prior works use diffusion models to generate driving images conditioned on the 3D object layout. However, those models are trained on small-scale datasets like nuScenes, which lack appearance and layout diversity. Moreover, overfitting often happens, where the trained models can only generate images based on the layout data from the validation set of the same dataset. In this work, we introduce a simulator-conditioned scene generation framework called SimGen that can learn to generate diverse driving scenes by mixing data from the simulator and the real world. It uses a novel cascade diffusion pipeline to address challenging sim-to-real gaps and multi-condition conflicts. A driving video dataset DIVA is collected to enhance the generative diversity of SimGen, which contains over 147.5 hours of real-world driving videos from 73 locations worldwide and simulated driving data from the MetaDrive simulator. SimGen achieves superior generation quality and diversity while preserving controllability based on the text prompt and the layout pulled from a simulator. We further demonstrate the improvements brought by SimGen for synthetic data augmentation on the BEV detection and segmentation task and showcase its capability in safety-critical data generation.
Figure 1: SimGen is a controllable scene generation paradigm conditioned on a simulator. It learns from real-world and simulated data and then can generate diverse driving scenes based on the simulator’s control conditions and text prompt.
[…]
@@ 1 Introduction
A high-quality and diverse training data corpus is crucial for autonomous driving research and development. However, it is costly and laborious to annotate the data. Synthetic data generation is a promising alternative to harvest annotated training data, which brings realistic images and notable performance improvements across tasks like object detection [10] and semantic segmentation [75]. Besides the realism of the generated images, there are two necessary conditions to consider for a practical synthetic data generator for autonomous driving: 1) Appearance diversity, which ensures the synthetic data can cover a spectrum of weather, environmental, and geographical conditions. 2) Layout diversity, namely the distribution of objects, should cover different traffic scenarios, including safety-critical situations that are rare to collect in the real world.
@@
We construct a dataset called DIVA to obtain the appearance and layout diversity of the training data. DIVA comprises two parts: the web data and the synthesized data from the simulator. On the one hand, web data covers a worldwide range of geography, weather, scenes, and traffic elements, preserving the appearance diversity of a wide range of traffic participants. We design a data curation pipeline to collect and label YouTube driving videos. On the other hand, virtual driving videos with the traffic flow replayed from trajectory datasets or generated by a safety-critical scenario generator [80] are collected from a driving simulator [36]. In short, DIVA dataset blends real-world appearances and virtual layouts, consisting of 147.5 hours of Diverse In-the-wild and Virtual driving datA.
@@ 2.2 DIVA-Sim: Layout Diversity from the Simulator
Figure 2: Constructing DIVA dataset. DIVA-Real (left) comprises driving videos collected from YouTube. We apply a Vision Language Model to filter out noisy images via a checklist and utilize off-the-shelf models to annotate text, depth, and semantic labels. Meanwhile, DIVA-Sim (right) employs scene records and control policies in a simulator to create map elements and objects. It can generate digital twins of real-world data and safety-critical scenes. Then various kinds of sensors placed in the simulation produce multimodal images. Ren.:rendered; T.D.: top-down view. Numbers and letters indicate the sequence of processes.
@@ 3 SimGen Framework
Figure 3: Illustration of SimGen. SimGen processes text and scene record as inputs. The text is feature-encoded and utilized in the subsequent modules, whereas the scene record undergoes a simulator rendering into simulated depth and segmentation (SimCond) and extra conditions (ExtraCond). SimCond, coupled with the text features, is fed into the CondDiff module that converts SimCond into RealCond, representing real depth and segmentation. Eventually, the text features, RealCond, and ExtraCond are inputted into the ImgDiff module, where an Adapter merges multi-source conditions into a unified control condition and generates driving scene images.
@@ 6 Conclusion
Acknowledgements
The project was supported by the NSF Grants CCRI-2235012 and RI-2339769, and the Sony Focused Research Award. YZ, HZ, and MG were supported by the National Natural Science Foundation of China (No. 62432008). ZP is supported by the Amazon Fellowship via the Science Hub for Humanity and Artificial Intelligence at UCLA.
@@ Appendix A Discussions
SimGen project page provides links to the YouTube videos (DIVA-Real_Video_Links) used in DIVA-Real, as well as digital twins of nuScenes dataset (DIVA-Sim_nuSc_Digital_Twins) and safety-critical video clips (DIVA-Sim_Safety-critical_Demo_Videos) included in DIVA-Sim.
@@
Q3. What is the definition of safety-critical scenarios and how to ensure they are realistic and feasible?
A safety-critical scenario is a situation where one or more vehicles collide with the ego vehicle, which is rare to collect in real-world datasets like Waymo. We utilize CAT [80] to generate risky behaviors from logged scenarios to ensure reality and feasibility, which uses a data-driven motion prediction model that predicts several modes of possible trajectories of each traffic vehicle. Please refer back to [80] for a detailed description of safety-critical scenarios.
Q4. Broader impact. What are potential applications and future directions with the provided DIVA data and the SimGen model, for both academia and industry?
@@
Q5. Limitations. What are the issues with current designs, and corresponding preliminary solutions?
@@
However, as a pioneering work exploring how to introduce simulators into generative models for diverse driving scene generation, the primary focus of this work is the simulator-conditioned generalization ability across unseen driving scenarios rather than multi-view designs and computational overhead. Future work might include trying cross-frame attentions [20, 38, 72, 14], faster sampling methods [9, 46, 63, 84], and transferring our general method to more efficient diffusion models.
@@ B.1.2 Multimodal Annotation
Table 7: Location distribution of nuScenes and DIVA-Real.
Dataset North America South America Europe Asian Africa
nuScenes 44.1% 0.0% 0.0% 55.9% 0.0%
DIVA-Real 56.9% 8.5% 16.9% 14.6% 3.1%
Table 8: Time period distribution of nuScenes and DIVA-Real.
Dataset Daytime Dawn Dusk Nighttime
nuScenes 88.4% 0.0% 0.0% 11.6%
DIVA-Real 55.8% 16.3% 10.1% 17.8%
Table 9: Weather distribution of nuScenes and DIVA-Real.
Dataset Normal Rainy Cloudy Foggy Snowy
nuScenes 80.5% 19.5% 0.0% 0.0% 0.0%
DIVA-Real 58.2% 1.0% 28.6% 2.1% 10.2%
@@ B.2.2 Layout Diversity Highlights
Table 10: Layout distribution of nuScenes and DIVA-Sim.
Dataset Forward Left Right Left Lane Right Lane Intersection U-Turn Stop
@@ B.4 License and Privacy Considerations
All the data is under the CC BY-NC-SA 4.0 licensehttps://creativecommons.org/licenses/by-nc-sa/4.0/deed.en. Other datasets (including nuScenes [53], Waymo Open [65], Metadrive [36]) inherit their own distribution licenses.
@@ C.2 Model Design
Realism-controllability trade-off. Apart from the discretization steps of the CondDiff solver, the critical hyperparameter for sim-to-real transformation is $t_{s}$, the starting time of the image synthesis process in the reverse SDE. We notice that with a fixed CondDiff model, there’s a trade-off between Realism and Controllability when choosing different $t_{s}$ values. Smaller $t_{s}$ values lead to fewer denoising steps, giving SimCond more control over image generation but potentially compromising realism. Generally, we find $t_{s}\in[0.4,0.65]$ to work well for the foreground, and we ultimately select $t_{s}$ as 0.5 for foregrounds.